# Making it to First: The Random Access Problem in DNA Storage

### v1

@@ 3 Optimal Random Access Expectation for Two Information Strands
In this section, we determine the exact values of $T_{q}(2)$ and $T(2)$. We start by identifying the structure of a $q$-ary two-row matrix $G$ that minimizes $T_{\max}(G)$.
@@
In this section, let $x_{1}$ and $x_{2}$ denote the number of columns in $G$ of the form $(1,0)^{T}$ and $(0,1)^{T}$ respectively (recall that these vectors correspond to the information strands). Observe that over ${\mathbb{F}}_{q}$ there are exactly $q-1$ vectors distinct from the information strands such that no two distinct vectors among them are collinear; the only columns we need to be concerned about are those that are not collinear because of Remark 1. Denoting by $\beta\in{\mathbb{F}}_{q}\setminus\{0\}$ a primitive element, these vectors can be expressed as $(1,\beta^{i})^{T}$ for $20\leq i\leq q-$ 2. Let $a_{i}$ denote the number of columns in $G$ of the form $(1,\beta^{i})^{T}$, then the total number of columns in $G$ is given by $x\triangleq x_{1}+x_{2}+\sum_{i=0}^{q-2}a_{i}$.
In the next claim we calculate the expectation of $\tau_{1}(G)$ and $\tau_{2}(G)$.
@@
It holds that:
@@
In the next lemma, we will demonstrate that it suffices to assume $x_{1}=x_{2}$, whenever we want to minimize $T_{\max}(G)$.
@@
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $x_{1}>x_{2}$, then there exists a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$.
@@
Assume that $x_{1}>x_{2}$. From Equation (3), we have
@@
Now, define the concatenated matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,0)^{T}$ is replaced by $(0,1)^{T}$, and vice versa.
For $G^{\prime}$, we have
@@
Since $x_{1}>x_{2}$, it follows that $T_{\max}(G^{\prime})<T_{\max}(G)$. This completes the proof of Lemma 2. ∎
The next lemma is analogous to Lemma 2 but for the $a_{i}$ values, i.e., the number of columns of the form $(1,\beta^{i})^{T}$ in $G$.
@@
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $a_{i}<a_{j}$ for some $i,j\in\{0,1,\dots,q-2\}$, then there exits a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$.
@@
Similarly to the proof of Lemma 2, define a matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,\alpha^{i})^{T}$ is replaced by $(1,\alpha^{j})^{T}$, and vice versa. Then we have
@@
By Lemmas 2 and 3 we can assume that in an optimal matrix $G$, $x_{1}=x_{2}$ and $a_{i}=a_{j}$ for all $i,j\in\{0,\dots,q-2\}$. Substituting these assumptions in Equation (3) (let $a:=a_{i}$ and so $x=2x_{1}+(q-1)a$), we obtain
@@
In the next theorem, we determine the value of $T_{q}(2)$.
@@
$a=\frac{\sqrt{2}q-2}{q^{2}-2}x_{1}.$ (7)
This value corresponds to a minimum for Equation (5), since the derivative is negative before this value and positive afterwards. Note that although the value of $a$ is not an integer, due to its continuous dependence on $x_{1}$, it is always possible to select $x_{1}$ such that the value of $a$ in Equation (7) is arbitrarily close to an integer.
Substituting the value of $a$ from Equation (7) into Equation (5), we find
@@
By taking $q$ to infinity we obtain the following corollary, which proves that the construction presented in [4, Theorem 12] is optimal.
@@
It holds that $T(2)=\liminf_{q\longrightarrow\infty}T_{q}(2)=1+\frac{2}{\sqrt{2}+1}\approx 0.914\cdot 2$.
4 General Construction
In this section, we construct a $k$-dimensional code $\mathcal{C}_{k}$, which generalizes the construction presented in [17] specific to $k=3$. The construction will be presented in terms of a generator matrix for $\mathcal{C}_{k}$. In our construction, all columns of the generator matrix of the code, which we denote by $G_{k}$, are of weight $1$ or $2$. Consequently, the columns of $G_{k}$ can be partitioned into $\binom{k}{2}+k$ disjoint sets, where each set of columns corresponds to either a vertex or an edge in $K_{k}$, the complete graph on $k$ vertices. More precisely, we label $K_{k}=(V_{k},E_{k})$ in the following way:
  • (i)
    the set of vertices $V_{k}$ of $K_{k}$ correspond to the $k$ basis vectors $\textbf{e}_{i}$, $i\in[k]$, and we label by $V_{i}$ the vertex corresponding to $\textbf{e}_{i}$;
  • (ii)
    an edge between $\textbf{e}_{i_{1}}$ and $\textbf{e}_{i_{2}}$ for $i_{1},i_{2}\in[k]$ corresponds to the set of columns of $G_{k}$ with support $\{i_{1},i_{2}\}$ and we denote by $E_{i,j}$ the set of columns with support $\{i,j\}$.
Our objective is to ensure that $G_{k}$ is recovery complete, which we formally define as follows:
@@ 4 General Construction
A matrix $G_{k}$ is called recovery complete if the following conditions hold:
@@
    For any $j_{1},j_{2}\in[k]$, given two distinct columns ${{\boldsymbol{g}}}_{i},{{\boldsymbol{g}}}_{j}\in E_{j_{1},j_{2}}$, it is possible to recover the information strands $j_{1}$ and $j_{2}$, i.e., $\{{{\boldsymbol{e}}}_{i},{{\boldsymbol{e}}}_{j}\}\subseteq\langle{{\boldsymbol{g}}}_{i},{{\boldsymbol{g}}}_{j}\rangle$.
@@
    Suppose we collect one column each from $E_{j_{1},j_{2}},E_{j_{2},j_{3}},\ldots E_{j_{m},j_{1}}$ where the edges $\{j_{1},j_{2}\},\ldots,\{j_{m},j_{1}\}$ form a cycle in $K_{k}$ of length $m\leq k$. Then, it is possible to recover the information strands corresponding to the indices ${j_{1}},\ldots,{j_{m}}$.
We now turn to constructing recovery complete matrices, and we will begin by revisiting the case where $k=3$. While a construction for this case, based on a geometric viewpoint, was presented in [17], the approach taken here is designed so that it can be applied to a more general framework. Our generator matrices for the case where $k=3$ will have the following form:
[…]
where $G_{3}(x)\in{\mathbb{F}}_{q}^{3\times(3+3x)}$ and $\beta$ is a primitive element of ${\mathbb{F}}_{q}$.
@@
$\begin{bmatrix}1&1\\ \beta^{i_{j}}&\beta^{i_{j^{\prime}}}\end{bmatrix}.$
Since the determinant of this matrix is non-zero whenever $i_{j}\not\equiv i_{j^{\prime}}\bmod(q-1)$, the result follows.
@@
The requirement in (12) is identical to the construction of sum-free sets [11, 15]; A sum-free set $\mathcal{S}$ is defined as a set for which there are no solutions to the equation $r+\ell=s$ with $r,\ell,s\in\mathcal{S}$. For the setup where the sum-free set is defined over an abelian group of order $n$ (in our case $n=q-1$), it is known that if $n\not\equiv 0\bmod 2$ (as is the case here, since $q$ is assumed to be a power of $2$), then there exists a sum-free set of size at least $n(1/3-\frac{1}{3n})$. Let $\mathcal{S}\subseteq\{0,\dots,q-2\}$ be a sum-free set of size $x$. Setting the powers of $\beta$ in $E_{1,2}$, $E_{1,3}$, and $E_{2,3}$ to be equal to the elements of $\mathcal{S}$, yields the desired result. Although the requirement $x\leq\frac{q-1}{2}$ in [17], which is optimal, is superior to ours, our construction still ensures $q=\mathcal{O}(x)$, demonstrating feasibility.
In Construction 1, we extend these ideas to the case where $k>3$.
@@
where there are $\binom{k}{2}$ sub-matrices $E_{i,j}$, each corresponding to a pair of distinct elements from $[k]$. Each submatrix $E_{i,j}$ has the following properties:
@@
Theorem 2 (see [5, Theorem 2]).
Let $q=2^{tk}$ for some $t\geq 1$. Then there exist $12^{t}+$ 1 integers $j_{0}=0,j_{1}=1,j_{2},\dots,j_{2^{t}+1}\leq q-2$ such that the sums
@@
For $q=2^{tk}$, where $t=\lceil\log_{2}(x\binom{k}{2}-1)\rceil$, there exists a set $\mathcal{S}$ as described in Construction 1, and the generator matrix $G_{k}(x)$ in (15) is recovery complete.
@@
We begin by constructing a set $\mathcal{S}$ as described in Construction 1. Let $t$ be the smallest integer such that $x\binom{k}{2}\leq 2^{t}+1$. From Theorem 2 there exists a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$ of size $12^{t}+$ 1 whenever $q=2^{tk}$, and we can take a subset $\mathcal{S}$ of this sequence of size $x\binom{k}{2}$ with $0\in\mathcal{S}$. Now if we have $k^{\prime}$ different elements $\{i_{1},\dots,i_{k^{\prime}}\}\in\mathcal{S}$ for which Equation (13) holds with equality, we can split the sum among the terms corresponding to positive $\ell_{i}$’s, and negative $\ell_{i}$’s, respectively, resulting in two different sums of (at most) $k-1$ terms which are equal mod $q-1$. Then we can add the appropriate number of 0’s to both sums (since $0\in\mathcal{S}$), so that both sums have exactly $k-1$ terms, and we get a contradiction since $\mathcal{S}$ is a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$.
Next, we show that $G_{k}(x)$ is recovery complete. The fact that two columns from $E_{i,j}$ have full rank follows from the same reasoning as in the case of $k=3$.
For the second condition of recovery completeness, consider the scenario where we select one column each from $E_{i_{1},i_{2}},E_{i_{2},i_{3}},$ $\dots,E_{i_{m},i_{1}},$ and where $\{i_{1},i_{2}\},\{i_{2},i_{3}\}$, $\ldots$, $\{i_{m},i_{1}\}$ is a cycle of length $m\leq k$. To satisfy the second property, we must show that the following matrix has full rank:
@@
where in every column $1<j<m$, one of the elements is $1$ and the other is a power of $\beta$ of the form $\beta^{i_{j}}$, $i_{j}\in\mathcal{S}$. Since our underlying field has characteristic $2$, the determinant of this matrix is given by:
@@
Let $T\subseteq\{2,3,\dots,m-1\}$ denote the subset such that $a_{i_{j}}=1$ for every $j\in T$ and $a_{i_{j}}=\beta^{i_{j}}$ for every $j\notin T$. Then the determinant can be expressed as:
@@
To ensure the determinant is non-zero, it suffices to require that:
$i_{m}+\sum_{j\notin T}i_{j}\not\equiv i_{1}+\sum_{j\in T}i_{j}\bmod q-1,$
where $i_{j}\neq i_{j^{\prime}}$ for $j\neq j^{\prime}$. This condition follows directly from property (13), ensuring that the matrix in (16) has full rank. Therefore, the generator matrix $G_{k}(x)$ from (15) is recovery complete. ∎
Observe that repeating $I_{k}$ multiple times in $G_{k}(x)$ does not affect the matrix’s recovery completeness. Henceforth, let $G_{k}(x,y)$ denote the matrix obtained from $G_{k}(x)$ by repeating $I_{k}$ $y$ times.
Before continuing with the analysis, we take a closer look at the structure of the recovery sets for recovery complete matrices. Similar to the discussion in Definition 1, we will find it useful to interpret the recovery process as one that is performed over the complete graph $K_{k}$ rather than over the columns of the matrix $G_{k}(x,y)$. More precisely, for every column, say $\bf g$, that is collected from $G_{k}(x,y)$, one of the following analogous events occurs over $K_{k}$:

### v2

@@ 3 Optimal Random Access Expectation for Two Information Strands
In this section, we determine the exact values of $T_{q}(2)$ and $T(2)$. To this end, we begin by analyzing the structure of the full-rank matrices that minimize $T_{q}(2)$, and in the process, we adopt the following notation.
@@
In this section, unless specified otherwise, $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix of full rank, and we let $x_{1}(G)$ and $x_{2}(G)$ denote the number of columns in $G$ that are equal to ${\boldsymbol{e}}_{1}=(1,0)^{T}$ and ${\boldsymbol{e}}_{2}=(0,1)^{T}$, respectively (recall that these vectors correspond to the information strands). Moreover, over ${\mathbb{F}}_{q}$, there are exactly q1q-1 vectors in ${\mathbb{F}}_{q}^{2}$ that are distinct from the information strands and mutually non-collinear. Due to Remark 1, we only need to consider such non-collinear columns for $G$. Denoting by $\beta\in{\mathbb{F}}_{q}\setminus\{0\}$ a primitive element of ${\mathbb{F}}_{q}$, the q1q-1 non-collinear vectors can be expressed as $(1,\beta^{i})^{T}$ for $0\leq i\leq q-2$. For $0\leq i\leq q-2$ let $a_{i}(G)$ denote the number of columns in $G$ of the form $(1,\beta^{i})^{T}$. The total number of columns in $G$ is then given by $x(G)\triangleq x_{1}(G)+x_{2}(G)+\sum_{i=0}^{q-2}a_{i}(G)$. When the matrix $G$ is clear from context, we omit specifying $G$ in our notation and simply write $x,x_{1},x_{2}$ and $a_{i}$ instead of $x(G),x_{1}(G),x_{2}(G)$ and $a_{i}(G)$.
In the next claim we express the expectations of $\tau_{1}(G)$ and $\tau_{2}(G)$ in terms of $x_{1}(G)$, $x_{2}(G)$ and the $a_{i}(G)$’s for $0\leq i\leq q-2$.
@@
For a matrix $G\in{\mathbb{F}}_{q}^{2\times n}$, it holds that:
@@
The following lemma shows that, in order to determine the value of $T_{q}(2)$, one can assume $x_{1}=x_{2}$.
@@
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $x_{1}(G)>x_{2}(G)$, then there exists a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$ and $x_{1}(G^{\prime})=x_{2}(G^{\prime})$.
@@
Denote $x_{1}=x_{1}(G),\text{\leavevmode\nobreak\ }x_{2}=x_{2}(G),\text{\leavevmode\nobreak\ }x=x(G),\text{\leavevmode\nobreak\ }a_{i}=a_{i}(G)$ for $0\leq i\leq q-2$, and assume that $x_{1}>x_{2}$. From Equation (3), we have
@@
Now, consider the concatenated matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,0)^{T}$ is replaced by $(0,1)^{T}$, and vice versa.
First, it is clear from the construction of $G^{\prime}$, that $x_{1}(G^{\prime})=x_{2}(G^{\prime})$. In addition, we have that:
@@
Since $x_{1}>x_{2}$, it follows that
$\frac{x_{1}+x_{2}}{2x-x_{1}-x_{2}}<\frac{x_{1}}{x-x_{1}},$
and thus $T_{\max}(G^{\prime})<T_{\max}(G)$. ∎
The next Lemma is the analogs of Lemma 2 for the $a_{i}^{\prime}s$.
@@
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $a_{i}<a_{j}$ for some $i,j\in\{0,1,\dots,q-2\}$, then there exists a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$.
@@
Similarly to the proof of Lemma 2, define the matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,\alpha^{i})^{T}$ is replaced by $(1,\alpha^{j})^{T}$, and vice versa. Then we have
@@
Note that, both Lemma 2 and Lemma 3 reduce not only the maximum random-access expectation but also the average. Specifically, in both lemmas it holds that $\frac{{\mathbb{E}}[\tau_{1}(G^{\prime})+{\mathbb{E}}[\tau_{2}(G^{\prime})]]}{2}\leq\frac{{\mathbb{E}}[\tau_{1}(G)+{\mathbb{E}}[\tau_{2}(G)]]}{2}.$
Now, by Lemmas 2 and 3, to determine the value of $T_{q}(2)$, it is sufficient to consider matrices $G$, for which $x_{1}=x_{2}$ and $a_{i}=a_{j}$ for all $i,j\in\{0,\dots,q-2\}$. Writing $x_{1}=x_{2}$, $a:=a_{i}$, and $x=2x_{1}+(q-1)a$, Equation (3) simplifies to
@@
The next theorem leverages this formula to determine the exact value of $T_{q}(2)$.
@@
$a^{*}=\frac{\sqrt{2}q-2}{q^{2}-2}x_{1}.$ (7)
This choice indeed minimizes the expression in Equation (5), since the the derivative is negative immediately before $a^{*}$ and positive immediately after it. Although $a^{*}$ is not generally an integer, its continuous dependence on $x_{1}$, allows us to select $x_{1}$ so that $a^{*}$ can be made arbitrarily close to an integer.
Substituting $a^{*}$ from Equation (7) into Equation (5) yields
@@
[…]
Figure 1: Normalized (by $k=2$) random access expectation $T_{q}(2)$ from Theorem 1 and various prime powers $q$.
We include a plot (see Figure 1) to show what the random access expectation from Theorem 1 is for various values of $q$. Letting $q\to\infty$ gives the following corollary, which confirms the optimality of the construction in [4, Theorem 12].
@@
It holds that $T(2)=\liminf_{q\to\infty}T_{q}(2)=1+\frac{2}{\sqrt{2}+1}\approx 0.914\cdot 2$.
4 Record-Low Random-Access Expectation
In this section, we will construct a rank-$k$ matrix, which we will denote by $G_{k}$, with a random access expectation that improves upon all previously known results for all $k\geq 4$. This construction is a generalization of the construction presented in [19], which was specifically given for $k=3$. In the construction, all columns of the matrix $G_{k}$ are of weight $1$ or $2$.
Accordingly, the columns of $G_{k}$ can be naturally interpreted as representing the vertices and edges of $K_{k}$, the complete graph on $k$ vertices in the following way: If we denote by $V_{i}$, $i\in[k]$ the vertices of $K_{k}$, then, for $i\in[k]$, the set of columns of weight $1$ with support $i$ correspond to the vertex $V_{i}$, and for $i,j\in[k]$ with $i\neq j$, the set of columns of weight $2$ with support $i,j$ correspond to the edge between $V_{i}$ and $V_{j}$. This perspective will prove useful in Section 4.2, where we analyze the random access expectation of the proposed construction.
In Section 4.1, we present the formal construction of $G_{k}$, but first we introduce an additional notation and definition. For $\in[k], iji\neq j$, let $E_{i,j}$ denote the set of columns of weight $2$ with support $\{i,j\}$ in $G_{k}$. Our objective is to ensure that $G_{k}$ satisfies a property we call recovery completeness, which we formally define below.
@@ 4 Record-Low Random-Access Expectation
A rank-$k$ matrix $G\in{\mathbb{F}}_{q}^{k\times n}$ is called recovery complete if the following conditions hold:
@@
    For any $i,j\in[k]$, given two distinct columns ${{\boldsymbol{g}}},{{\boldsymbol{g}}}^{\prime}\in E_{i,j}$, it is possible to recover the basis vectors ${\boldsymbol{e}}_{i}$} and ${\boldsymbol{e}}_{j}$, i.e., $\{{{\boldsymbol{e}}}_{i},{{\boldsymbol{e}}}_{j}\}\subseteq\langle{{\boldsymbol{g}}},{{\boldsymbol{g}}}^{\prime}\rangle$.
@@
    Suppose we collect one column each from $E_{j_{1},j_{2}},E_{j_{2},j_{3}},\ldots,E_{j_{m},j_{1}}$ so all the $j_{i}$’s are different and the edges $\{j_{1},j_{2}\},\ldots,\{j_{m},j_{1}\}$ form a cycle in $K_{k}$ of length $m\leq k$ . Then, it is possible to recover the basis vectors corresponding to the indices ${j_{1}},\ldots,{j_{m}}$.
As we will show later in Secion 4.2, recovery complete matrices have a good random access expectation. But first, we construct such matrices in the next section.
4.1 Construction of Recovery Complete Matrices
We begin by revisiting the case where $k=3$. While a construction for this case, based on a geometric viewpoint, was presented in [19], the approach taken here is designed so that it can be applied to a more general framework. Our matrices for the case where $k=3$ will have the following form:
$G_{3}(x)=\left[\begin{array}[]{ccc|cccc|cccc|cccc}1&0&0&1&1&\cdots&1&1&1&\cdots&1&0&0&\cdots&0\\ 0&1&0&\beta^{i_{1}}&\beta^{i_{2}}&\cdots&\beta^{i_{x}}&0&0&\cdots&0&1&1&\cdots&1\\ 0&0&1&0&0&\cdots&0&\beta^{i_{x+1}}&\beta^{i_{x+2}}&\cdots&\beta^{i_{2x}}&\beta^{i_{2x+1}}&\beta^{i_{2x+2}}&\cdots&\beta^{i_{3x}}\end{array}\right]$ (11)
$\hskip 93.89418pt\underbrace{\hskip 93.89418pt}_{E_{1,2}}\hskip 5.69046pt\underbrace{\hskip 113.81102pt}_{E_{1,3}}\hskip 5.69046pt\underbrace{\hskip 125.19194pt}_{E_{2,3}}$
where $G_{3}(x)\in{\mathbb{F}}_{q}^{3\times(3+3x)}$, $\beta$ is a primitive element of ${\mathbb{F}}_{q}$ and all the $i_{j}$’s are different.
@@ 4.1 Construction of Recovery Complete Matrices
Claim 2.
Let $\{i_{1},\dots,i_{3x}\}\subseteq\{0,\dots,q-2\}$ be a set of distinct integers such that for all distinct $s,r,\ell\in\{i_{1},\dots,i_{3x}\}$, we have
$s\not\equiv r+\ell\bmod(q-1).$
Then, the matrix $G_{3}(x)$, as described in (11), is recovery complete.
Proof.
@@
$\begin{bmatrix}1&1\\ \beta^{r}&\beta^{\ell}\end{bmatrix}$
for some $r,\ell\in\{i_{1},\dots,i_{3x}\}$. Since the determinant of this matrix is non-zero whenever $r\not\equiv\ell\bmod(q-1)$, and we assumed that all the powers of $\beta$ showing up in $G_{3}(x)$ are different and in $\{0,\dots,q-2\}$, the first property of recovery completeness trivially follows.
@@
By the assumptions in the claim, the condition in (12) holds. ∎
The requirement in (12) is identical to the construction of sum-free sets [13, 17]; A sum-free set $\mathcal{S}$ is a set for which there are no solutions to the equation $r+\ell=s$ with $r,\ell,s\in\mathcal{S}$. For the setup where the sum-free set is defined over an abelian group of order $n$ (in our case n=q1n=q-1), it is known that if $n\not\equiv 0\bmod 2$ (as is the case here, since $q$ is assumed to be a power of $2$), then there exists a sum-free set of size at least $n(1/3-\frac{1}{3n})$.
Let $\mathcal{S}\subseteq\{0,\dots,q-2\}$ be a sum-free set of size $3x$. Setting the powers of $\beta$ in $E_{1,2}$, $E_{1,3}$, and $E_{2,3}$ to be equal to the elements of $\mathcal{S}$, yields the desired result. Although the requirement $x\leq\frac{q-1}{2}$ in [19], which is optimal, is superior to ours, our construction still ensures $q=\mathcal{O}(x)$, and can be naturally extended to larger values of $k>3$, as demonstrated in Construction 1.
@@
where there are $\binom{k}{2}$ sub-matrices $E_{i,j}\in{\mathbb{F}}_{q}^{k\times x}$, each corresponding to some $i,j\in[k]$ with $i\neq j$. Each submatrix $E_{i,j}$ has the following properties:
@@
Theorem 2 (see [7, Theorem 2]).
Let $q=2^{tk}$ for some $t\geq 1$. Then, there exist $2^{t}+1$ integers $j_{0}=0,j_{1}=1,j_{2},\dots,j_{2^{t}+1}\leq q-2$ such that the sums
@@
Let $q=2^{tk}$, where $t=\lceil\log_{2}(x\binom{k}{2}-1)\rceil$. Then, there exists a set $\mathcal{S}$ as described in Construction 1, and the generator matrix $G_{k}(x)$ in (15) is recovery complete.
@@
We begin by constructing a set $\mathcal{S}$ as described in Construction 1. Let $t$ be the smallest integer such that $x\binom{k}{2}\leq 2^{t}+1$. From Theorem 2 there exists a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$ of size $2^{t}+1$ whenever $q=2^{tk}$. We can then choose a subset $\mathcal{S}$ of this sequence of size $x\binom{k}{2}$ with $0\in\mathcal{S}$.
Assume toward contradiction, that there exist $k^{\prime}$ different elements $\{i_{1},\dots,i_{k^{\prime}}\}\in\mathcal{S}$ for which Equation (13) holds with equality. We can then split the sum among the terms corresponding to positive $\ell_{i}$’s, and negative $\ell_{i}$’s, respectively, resulting in two different sums of (at most) k1k-1 terms which are equal mod q1q-1. As $0\in\mathcal{S}$, we can pad both sums with zeroes to reach exactly k1k-1 terms. This contradicts the assumption that $\mathcal{S}$ is a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$.
We now show that $G_{k}(x)$ is recovery complete. The fact that two distinct columns from $E_{i,j}$ are linearly independent follows from the same reasoning as in the case of $k=3$.
To prove the second condition of recovery completeness, consider the scenario where we draw one column each from $E_{i_{1},i_{2}},E_{i_{2},i_{3}},$ $\dots,E_{i_{m},i_{1}},$ and where $\{i_{1},i_{2}\},\{i_{2},i_{3}\}$, $\ldots$, $\{i_{m},i_{1}\}$ is a cycle of length $m\leq k$. To satisfy the second property, we have to show that the matrix made of the corresponding columns of $G_{k}(x)$ is full rank, i.e., has rank $m$. In order to do so, note that we can delete all 0-rows, and permute rows in a suitable way, so that we end up with a matrix of the following form:
@@
where in every column $1<j<m$, $\{a_{i_{j}},b_{i_{j}}\}=\{1,\beta^{i_{j}}\}$, i.e., one of the non-zero entries is $1$ and the other is a power of $\beta$ of the form $\beta^{i_{j}}$, where $i_{j}\in\mathcal{S}$. Since our underlying field has characteristic $2$, the determinant of this matrix is given by:
@@
Let $T\subseteq\{2,3,\dots,m-1\}$ denote the subset such that $a_{i_{j}}=1$ for every $j\in T$, and $a_{i_{j}}=\beta^{i_{j}}$ for every $j\notin T$. Then the determinant becomes:
@@
To ensure that this expression is nonzero, it suffices to require:
$i_{m}+\sum_{j\notin T}i_{j}\not\equiv i_{1}+\sum_{j\in T}i_{j}\bmod(q-1),$
where all $i_{j}$ are distinct. This condition follows directly from equation (13), which is satisfied by the set $S$. Therefore, the matrix in (16) has full rank, completing the proof that the generator matrix $G_{k}(x)$ from (15) is recovery complete. ∎
In the construction described thus far, each information strand (i.e., column of weight 1) appears exactly once in $G_{k}(x)$. However, by repeating these columns multiple times, we can reduce the random access expectation. Importantly, such repetition does not affect the recovery completeness of the matrix. Henceforth, let $G_{k}(x,y)$ denote the matrix obtained from $G_{k}(x)$ by appending y1y-1 copies of the identity matrix $I_{k}$ to $G_{k}(x)$, resulting in a matrix with $y$ copies of each column of weight 1.
In the next section, we will evaluate the asymptotic random access expectation of $G_{k}(x,y)$ as $x,y\to{}\infty$, which will enable us to derive an upper bound on $T(k)$.
4.2 The Random Access Expectation of the Construction
Before analyzing the random access expectation of $G_{k}(x,y)$, we will find it useful to interpret the recovery process as one that is performed over the complete graph $K_{k}$ (see the beginning of Section 4) rather than over the columns of the matrix $G_{k}(x,y)$. More precisely, for each collected column $\bf g$ from $G_{k}(x,y)$, one of the following corresponding events occurs in $K_{k}$: